# SpIDER: Spatially Informed Dense Embedding Retrieval for Software Issue Localization

### v1

@@ A.3 Additional Experiments
Table 9: LocAgent performance on SWEBench-Verified and the Python-only subset of SWE-PolyBench. Running LocAgent took over 48 hours on SWEBench-Verified and approximately 24 hours on SWE-PolyBench, highlighting a key limitation of this approach. In addition to the long runtime, LocAgent also requires significantly more LLM invocations per instance compared to dense embedding retrieval and BM25, resulting in higher computational costs.
@@
Table 10: Comparison of LocAgent with dense retrieval methods + reranker for Python repositories in SWE-PolyBench and SWEBench-Verified
@@
SWE-PolyBench LocAgent 0.23 0.17 0.21
SweRankEmbed-Small ZS+ SpIDER 0.61 0.56 0.63
+ Reranker
CodeSAGE-Small ZS + SpIDER 0.– 0.– 0.–
+ Reranker
CodeSAGE-Small-PEFT + SpIDER 0.52 0.46 0.56
+ Reranker
BM25+ SpISR 0.– 0.– 0.–
+ Reranker
@@
Table 12: SpIDER + SweRankEmbed-Small performance with varying neighbor filtering threshold NN for $K=20$, $C=3$, $d=4$. Increasing NN improves coverage and Recall@20 with modest increases in LLM token usage. Avg. tokens indicate average consumption per LLM call; number of calls equals $C$. Best Recall and Accuracy per language are highlighted.

### v2

@@ A.10 Additional Experiments
Table 8: LocAgent performance on SWEBench-Verified and a subset of Python instances in SWE-PolyBench. Running LocAgent took over 48 hours on SWEBench-Verified and approximately 24 hours on SWE-PolyBench, highlighting a key limitation of this approach. In addition to the long runtime, LocAgent also requires significantly more LLM invocations per instance compared to dense embedding retrieval and BM25, resulting in higher computational costs.
@@
Table 9: Comparison of LocAgent with dense retrieval methods + LLM reranker for Python repositories in SWE-PolyBench
@@
Table 11: Ablation on primary neighborhood filtering threshold NN using SpIDER + SweRankEmbed-Small for $K=20$, $C=3$, $d=4$. Increasing NN improves coverage and Recall@20 with modest increases in LLM token usage. Avg. tokens indicate average consumption per LLM call; number of calls equals $C$. Best Recall and Accuracy per language are highlighted.
@@
Table 14: Comparison of dense retrieval and dense retrieval + LLM reranker for $K=20$ and $N=500$, $C=5$, $d=4$ on SWE-PolyBench. Winners per metric per embedding model are highlighted in light blue.
Language Model Retrieval Dense Retrieval Only Dense + LLM reranker
Recall@20 Accuracy@20 MRR@20 Recall@3 Accuracy@3 MRR@3
Python SweRankEmbed-Small ZS SpIDER 0.49 0.40 0.30 0.44 0.36 0.54
DER 0.42 0.35 0.26 0.38 0.31 0.48
CodeSAGE-Small ZS SpIDER 0.27 0.21 0.13 0.23 0.18 0.31
DER 0.21 0.17 0.11 0.19 0.16 0.25
CodeSAGE-Small-PEFT SpIDER 0.37 0.27 0.24 0.32 0.25 0.42
DER 0.29 0.20 0.21 0.26 0.20 0.38
BM25 SpISR 0.33 0.26 0.19 0.30 0.23 0.38
SR 0.31 0.24 0.18 0.27 0.22 0.35
Java SweRankEmbed-Small ZS SpIDER 0.36 0.27 0.18 0.25 0.17 0.36
DER 0.31 0.24 0.18 0.24 0.16 0.33
CodeSAGE-Small ZS SpIDER 0.23 0.17 0.10 0.16 0.09 0.26
DER 0.15 0.10 0.07 0.11 0.07 0.19
CodeSAGE-Small-PEFT SpIDER 0.28 0.18 0.15 0.20 0.14 0.28
DER 0.21 0.14 0.13 0.17 0.11 0.25
BM25 SpISR 0.16 0.10 0.10 0.12 0.08 0.20
SR 0.15 0.10 0.10 0.12 0.09 0.19
DER 0.52 0.45 0.16 0.41 0.34 0.42
CodeSAGE-Small ZS SpIDER 0.38 0.32 0.09 0.30 0.25 0.31
DER 0.26 0.21 0.07 0.22 0.17 0.24
CodeSAGE-Small-PEFT SpIDER 0.55 0.48 0.15 0.41 0.35 0.41
DER 0.45 0.38 0.14 0.36 0.30 0.39
BM25 SpISR 0.46 0.39 0.13 0.36 0.30 0.38
SR 0.36 0.29 0.11 0.31 0.25 0.34
DER 0.29 0.21 0.18 0.25 0.19 0.30
CodeSAGE-Small ZS SpIDER 0.16 0.10 0.03 0.14 0.09 0.19
DER 0.10 0.06 0.06 0.09 0.06 0.14
CodeSAGE-Small-PEFT SpIDER 0.25 0.18 0.04 0.22 0.16 0.27
DER 0.19 0.13 0.11 0.18 0.12 0.24
BM25 SpISR 0.11 0.09 0.02 0.10 0.07 0.11
SR 0.07 0.05 0.04 0.07 0.05 0.09
Table 15: Dense retrieval performance across benchmarks for $K=20$, $N=500$, $C=5$, $d=4$. Winners per metric per embedding model highlighted in light blue.
Language Model Retrieval Recall@20 Accuracy@20 MRR@20
Multi-SWEBench
Java SweRankEmbed-Small ZS SpIDER 0.37 0.28 0.08
DER 0.32 0.24 0.08
CodeSAGE-Small ZS SpIDER 0.21 0.19 0.05
DER 0.15 0.13 0.03
CodeSAGE-Small-PEFT SpIDER 0.31 0.25 0.10
DER 0.23 0.17 0.09
BM25 SpISR 0.23 0.17 0.07
SR 0.22 0.17 0.05
JavaScript SweRankEmbed-Small ZS SpIDER 0.39 0.30 0.18
DER 0.31 0.23 0.15
CodeSAGE-Small ZS SpIDER 0.17 0.12 0.08
DER 0.09 0.06 0.05
CodeSAGE-Small-PEFT SpIDER 0.32 0.24 0.13
DER 0.25 0.18 0.11
BM25 SpISR 0.21 0.14 0.08
SR 0.14 0.10 0.06
TypeScript SweRankEmbed-Small ZS SpIDER 0.52 0.42 0.08
DER 0.40 0.32 0.23
CodeSAGE-Small ZS SpIDER 0.33 0.26 0.04
DER 0.26 0.21 0.13
CodeSAGE-Small-PEFT SpIDER 0.30 0.26 0.05
DER 0.19 0.15 0.12
BM25 SpISR 0.32 0.27 0.04
SR 0.24 0.21 0.11
SWEBench-Verified
Python SweRankEmbed-Small ZS SpIDER 0.61 0.56 0.32
DER 0.54 0.49 0.29
CodeSAGE-Small ZS SpIDER 0.30 0.27 0.13
DER 0.21 0.19 0.10
CodeSAGE-Small-PEFT SpIDER 0.55 0.49 0.27
DER 0.43 0.38 0.22
BM25 SpISR 0.45 0.41 0.25
SR 0.38 0.34 0.22
$\boldsymbol{C}$ Python JavaScript Java TypeScript
1 0.47 0.38 0.29 0.56 0.48 0.18 0.32 0.24 0.18 0.32 0.24 0.06
3 0.49 0.39 0.30 0.59 0.51 0.18 0.35 0.26 0.19 0.31 0.24 0.06
5 0.49 0.40 0.30 0.60 0.52 0.18 0.36 0.27 0.18 0.35 0.28 0.06
7 0.49 0.39 0.30 0.60 0.52 0.18 0.36 0.27 0.19 0.37 0.30 0.06
9 0.48 0.39 0.30 0.60 0.53 0.19 0.37 0.27 0.19 0.39 0.31 0.06
MRR@$K$ Gains Depend on Graph Neighborhoods